OpenStudy (anonymous):

Can someone please solve and show all your steps? Thanks! 2/3 + y = 3/4

OpenStudy (anonymous):

You have to make the bases equal. The equation becomes 8/12 + y = 9/12 y= 1/12 If that makes sense... :P